8z+9=5 what does z equal?
Answer by hmj(6)

First try to get 8z on one side, so subtract the 9. In you do something to one side then you have to do it to the other, so
8z+9=5 (subtracting 9 from both sides gives you...)
8z=-4
now you have to divide both side by 8 to find the value of z
so,
z=-.5
To check your work, replace z with -.5
8(-.5)+9=5
-4+9=5
5=5
Since that works out, the answer is
z=(-.5)